OpenStudy (anonymous):

A plane has 270 gallons of fuel. The plane averages 2 2/9 miles per gallon. How far can the plane go?

OpenStudy (anonymous):

600 miles

OpenStudy (anonymous):

270/(2 2/9) =270/ =600

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

\[total distance covered = 270 gallons \times \frac{20}{9} miles/gallon = 30 \times 20 miles = 600 miles\]

OpenStudy (anonymous):

\[20/9 * 270/1 = 5400/9\therefore Distance=600 miles\]